what volume of HCL 0.2 molar should add to 250 mlit sodium mandelat 0.3 and finally they made a bafer soloution with PH 3.37
1
Expert's answer
2012-11-06T11:28:45-0500

pH = pKa + lg (C salt * V salt / C acid * V acid)
3.37 = 3.41 + lg (0.25*0.3 / 0.2*V acid)
-0.04 = lg (0.25*0.3 / 0.2*V acid)
0.25*0.3 / 0.2*V acid = 0.912
V acid = 0.411 L
V acid = 411 ml